OpenStudy (anonymous):

for adding/subtracting complex numbers, if this "(8+4i)-(3+2i)" is the question, would the answer be 5+6i? for adding/subtracting complex numbers, if this "(8+4i)-(3+2i)" is the question, would the answer be 5+6i? @Mathematics

OpenStudy (anonymous):

no thats incorrect. you need to distribute the negative sign

OpenStudy (anonymous):

I would do it this way (8 + 4i) - (3 +2i) = 8 - 3 + 4i - 2i =5 + 2i